On taking 60.0g CH3COOH and 46.0g CH3CH2OH in a 5L flask in the presence of H3O+(catalyst), at 298 K 44.0g of CH3COOC2H5 is formed at equilibrium. If amount of CH3COOH is doubled without affecting the amount of CH3CH2OH, then CH3COOC2H5 formed is

Grade:12

1 Answers

Vikas TU
7 years ago
After doubling the amount,
Moles of CH3COOH => 120/60 = 2 moles.
Moles of CH3CH2OH => 46/46 = 1 moles.
POAC on Carbon,
2 * 2 + 1 * 2 = 2*n
we get,
n = 6/2 = 3 moles.
Mass of CH3COOC2H5 => 3 * 88 = >  264 gm.